# SignSpeak: Open-Source Time Series Classification for ASL Translation

### v1

@@ 1 Introduction
Sensor-based models using embedded devices have been introduced to treat ASL as a time-series multi-label classification problem to address the limitations of optical systems. However, many of these datasets are private [6, 7] and have not been trained on a well-practiced sign-based language such as ASL [8], limiting their applicability. To address this, we introduce SignSpeak, an open-source ASL dataset comprising of 7200 recordings of 36 classes. Our dataset was recorded using five low-cost flex sensors, one for each finger, for all letters and numbers in ASL. The scale of our dataset enables researchers to test novel models on a dataset collected in an environment aimed at real-world feasibility to progress ASL-to-speech efforts. We extensively benchmarked various methods on this dataset, with our best result achieving 92% categorical accuracy, which matches or exceeds previous ASL time-series classification work [9].111 The GitHub codebase and dataset are available at https://github.com/adityamakkar000/ASL-Sign-Research and https://doi.org/10.7910/DVN/ODY7GH respectively.

### v2

@@ 1 Introduction
Sensor-based models using embedded devices have been introduced to treat ASL as a time-series multi-label classification problem to address the limitations of optical systems. However, many of these datasets are private [6, 7] and have not been trained on a well-practiced sign-based language such as ASL [8], limiting their applicability. To address this, we introduce SignSpeak, an open-source ASL dataset comprising of 7200 recordings of 36 classes. Our dataset was recorded using five low-cost flex sensors, one for each finger, for all letters and numbers in ASL. The scale of our dataset enables researchers to test novel models on a dataset collected in an environment aimed at real-world feasibility to progress ASL-to-speech efforts. We extensively benchmarked various methods on this dataset, with our best result achieving 92% categorical accuracy, which matches or exceeds previous ASL time-series classification work [9].111 The GitHub codebase and dataset are available at
https://github.com/adityamakkar000/ASL-Sign-Research
and https://doi.org/10.7910/DVN/ODY7GH respectively.